OpenStudy (anonymous):

Who can prove this simply? 2x2 = 4

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

2+2=4

OpenStudy (anonymous):

\[2^2\] <--- simplified

OpenStudy (anonymous):

but I think 2^2 = 8 and 2x2 = 8!!! prove that i'm wrong...

OpenStudy (anonymous):

2 squared is 2x2...

OpenStudy (anonymous):



OpenStudy (anonymous):

ok,it's good :) anybody else can prove it?

OpenStudy (anonymous):

If we assume that 2+2 = 4 and that x*2 = x+x then 2*2=4.

ganeshie8 (ganeshie8):

By definition of multiplication \(\color{purple}{a} \times \color{Red}{b}\) is, the quantity \(\color{Red}{b}\) taken \(\color{purple}{a}\) times \(\color{purple}{a} \times \color{Red}{b} = \color{Red}{b} + \color{Red}{b} + ..... \color{purple}{a}\) times By definition of multiplication \(\color{purple}{2} \times \color{Red}{2}\) is, the quantity \(\color{Red}{2}\) taken \(\color{purple}{2}\) times \( \color{purple}{2} \times \color{Red}{2} = \color{Red}{2} + \color{Red}{2} + ..... \color{purple}{2}\) times

ganeshie8 (ganeshie8):

oh that just 2 + 2 which is 4

ganeshie8 (ganeshie8):

so 2x2 = 2 + 2 + .... 2 times = 4

OpenStudy (anonymous):

yes @plitter and @ganeshie8 proved it very simply :)

OpenStudy (anonymous):

@paigeRG , your work was great too :)

OpenStudy (anonymous):

Yaaaayyy! :D

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

but i had another way to prove it... how about using "Proof by contradiction" ?

ganeshie8 (ganeshie8):

oh you're not going to ask why 2+2 = 4... :)

ganeshie8 (ganeshie8):

we can try...

OpenStudy (anonymous):

Let's assume that 2*2 is not 4 then 2+2 is not 4 either. Which it is! Thus 2*2 has to be 4. QED!

OpenStudy (anonymous):